# UMGAD: Unsupervised Multiplex Graph Anomaly Detection

### v2

@@ V-B Anomaly Score Threshold Selection for Unsupervised Anomaly Detection (RQ1)
The results on four datasets are shown in Fig. 2, which shows the trend of the ranked node anomaly scores for the best-performing baselines, TAM, ADA-GAD, GADAM, AnomMAN, and for our UMGAD, respectively. We can see that compared with the other four best-performing baselines, the curve of our UMGAD can converge quickly (become stable) to a position closer to the actual number of anomalies in the datasets. It allows UMGAD to distinguish better the anomaly scores of anomalous nodes from those of normal nodes, and thus be able to identify anomalies more accurately. Notably, a robust anomaly detection model should effectively separate anomalous nodes from normal ones by leveraging anomaly scores. Specifically, the anomaly scores, when arranged in descending order, ought to display a distinct inflection point where they stabilize. Nodes with higher anomaly scores preceding this inflection point are likely to be anomalous, whereas those with lower, more uniform scores beyond this point can be viewed as normal. Therefore, we use the anomaly score at inflection point of the fitting curve as the anomaly threshold. The number of anomalous nodes corresponding to the selected threshold is very close to the real number of anomalies which can be identified by ground truth information, which proves that our method can effectively alleviate the difficult problem of selecting anomaly thresholds.
Our proposed anomaly threshold selection strategy is the first truly unsupervised method. It does not require any ground-truth information from the test set, making it a fair and objective model evaluation approach. Therefore, the experimental results in Table II are objective. Although the results in Table IV are generally better than those in Table II, they are not objective, as ground-truth information from the test set (e.g., the number of anomalous nodes) is used when selecting anomaly thresholds.
@@ V-C Performance Comparison in the Real Unsupervised Scenario (RQ2)
Next, we evaluate the performance of our UMGAD and all baselines in real unsupervised scenarios. The experimental results are shown in Table II. The best results are highlighted in bold, and the second-best results are underlined.
As we can see, our UMGAD obtains the optimal performance, significantly achieves 13.48%, and 11.68% average improvement in terms of AUC and macro-F1 across four datasets. We can draw the following conclusions: First, according to our proposed threshold selection strategy and the node anomaly score curves in Fig. 2, the number of detected anomalies corresponding to our selected thresholds is the closest to the real situation compared to other methods. Second, our UMGAD method considers the multi-relational correlation between nodes, and collaboratively combines GAEs with various masking mechanisms across different types of relations, effectively capturing diverse anomaly signals in both original and augmented views. Therefore, the anomaly scores of anomalous nodes can be better distinguished from those of normal nodes.
TABLE III: The comparison of UMGAD and its variants on AUC and macro-F1 (F1 for short) metrics.
@@
w/o M 0.712 0.669 0.783 0.689 0.829 0.716 0.571 0.556
w/o O 0.720 0.682 0.788 0.696 0.835 0.727 0.577 0.565
w/o A 0.714 0.670 0.782 0.691 0.827 0.719 0.571 0.558
w/o NA 0.731 0.690 0.795 0.706 0.842 0.735 0.582 0.575
w/o SA 0.726 0.685 0.788 0.701 0.839 0.730 0.580 0.567
w/o DCL 0.748 0.702 0.806 0.718 0.848 0.755 0.588 0.577
UMGAD 0.770 0.722 0.825 0.740 0.878 0.779 0.608 0.597
V-D Ablation Study (RQ3)
@@
Table III presents the results of model variants on four datasets in real unsupervised scenarios. All five variants underperform UMGAD in terms of AUC and macro-F1, underscoring the necessity and effectiveness of the proposed components. Among them, w/o M performs the worst, highlighting the importance of the graph-masked autoencoder (GMAE) in GAD tasks. Ordinary GAE struggles to accurately reconstruct the attributes and structural features of anomalous nodes, leading to suboptimal anomaly detection performance. w/o O outperforms w/o M and w/o A, emphasizing the critical role of the original multiplex heterogeneous graph. Relying solely on the augmented graph risks losing crucial anomaly information. The results for w/o NA and w/o SA further demonstrate that the augmented-view GMAE is effective in capturing anomalous information. Changes to node attributes and structural features in the augmented graph help mitigate the homogeneity trap common in GNN-based autoencoders. Finally, while w/o CDL performs the best among all variants, it still falls short of UMGAD, confirming the significant contribution of dual-view contrastive learning in optimizing node representations.
@@
Figure 4: The effect of masking ratio and masking subgraph size. The X-axis indicates the mask ratio, and 4 to 16 in the legend indicates the subgraph size.
V-E Parameter Sensitivity Analysis (RQ4)
V-E1 The effect of hyperparameters $\lambda$, $\mu$ and $\Theta$ in the final loss function
$\lambda$, $\mu$ and $\Theta$ correspond to the importance parameters of the two enhanced views and dual-view contrastive learning, respectively. Considering node attribute-level and subgraph-level enhanced views on top of the original view can improve the model performance, and the results of the parameter sensitivity experiments are shown in Fig. 3. The model performance is optimal when $\lambda$ and $\mu$ are set to 0.3 and 0.3, 0.3 and 0.4, 0.4 and 0.4, and 0.4 and 0.5 on the four datasets, respectively. Besides, our UMGAD tends to perform well by setting $\Theta$ to 0.1 across all datasets.
V-E2 The effect of masking ratio $r_{m}$ and masking subgraph size $|\mathcal{V}_{m}|$
@@ V-E1 The effect of hyperparameters $\lambda$, $\mu$ and $\Theta$ in the final loss function
Figure 5: The effect of weights $\alpha$ and $\beta$ in four datasets. The corresponding AUC values for the Alibaba and Amazon datasets are on the left Y-axis, and the corresponding AUC values for Retail and YelpChi are on the right Y-axis.
V-E3 The effect of weights $\alpha$ and $\beta$ that balance the importance of attribute and structure reconstruction
We conduct a sensitivity analysis of $\alpha$ and $\beta$, which balance attribute and structure reconstructions in the original and augmented views. As shown in Fig. 5, the model’s performance declines sharply for extreme values of $\alpha$ and $\beta$ $(<0.2\text{ or }>0.8)$, highlighting the importance of maintaining a balance. The AUC first increases and then decreases as $\alpha$ and $\beta$ grow, with optimal performance of our proposed UMGAD at $\alpha=0.5,0.5,0.6,0.5$ and $\beta=0.4,0.4,0.3,0.3$ across the four datasets.
TABLE IV: Performance comparison of all models on four datasets in the scenario of anomaly score threshold selection with ground truth leakage.
@@
Trad. Radar [IJCAI’17] 0.678$\pm$0.003 0.570$\pm$0.001 0.695$\pm$0.011 0.621$\pm$0.009 0.659$\pm$0.003 0.564$\pm$0.002 0.515$\pm$0.003 0.502$\pm$0.005
MPI ComGA [WSDM’22] 0.702$\pm$0.009 0.648$\pm$0.003 0.761$\pm$0.008 0.652$\pm$0.002 0.715$\pm$0.004 0.630$\pm$0.007 0.543$\pm$0.011 0.497$\pm$0.006
RAND [ICDM’23] 0.726$\pm$0.006 0.678$\pm$0.007 0.784$\pm$0.008 0.677$\pm$0.004 0.730$\pm$0.013 0.645$\pm$0.006 0.554$\pm$0.007 0.504$\pm$0.008
TAM [NeurIPS’24] 0.753$\pm$0.001 0.703$\pm$0.001 0.795$\pm$0.001 0.715$\pm$0.001 0.840$\pm$0.001 0.732$\pm$0.001 0.582$\pm$0.001 0.554$\pm$0.001
CL CoLA [TNNLS’21] 0.629$\pm$0.009 0.559$\pm$0.006 0.647$\pm$0.014 0.613$\pm$0.003 0.661$\pm$0.006 0.579$\pm$0.002 0.480$\pm$0.005 0.489$\pm$0.009
ANEMONE [CIKM’21] 0.689$\pm$0.004 0.631$\pm$0.004 0.710$\pm$0.007 0.632$\pm$0.009 0.704$\pm$0.002 0.614$\pm$0.005 0.541$\pm$0.005 0.511$\pm$0.003
Sub-CR [IJCAI’22] 0.704$\pm$0.009 0.624$\pm$0.008 0.728$\pm$0.006 0.648$\pm$0.008 0.675$\pm$0.018 0.591$\pm$0.003 0.520$\pm$0.014 0.491$\pm$0.008
ARISE [TNNLS’23] 0.740$\pm$0.018 0.683$\pm$0.005 0.774$\pm$0.006 0.669$\pm$0.003 0.768$\pm$0.006 0.679$\pm$0.002 0.565$\pm$0.001 0.525$\pm$0.007
SL-GAD [TKDE’21] 0.725$\pm$0.003 0.667$\pm$0.007 0.778$\pm$0.005 0.672$\pm$0.003 0.725$\pm$0.009 0.642$\pm$0.010 0.553$\pm$0.011 0.504$\pm$0.005
PREM [ICDM’23] 0.724$\pm$0.001 0.675$\pm$0.003 0.782$\pm$0.011 0.670$\pm$0.009 0.735$\pm$0.007 0.649$\pm$0.006 0.557$\pm$0.003 0.513$\pm$0.007
GCCAD [TKDE’22] 0.731$\pm$0.003 0.677$\pm$0.005 0.788$\pm$0.002 0.689$\pm$0.003 0.733$\pm$0.001 0.647$\pm$0.005 0.553$\pm$0.002 0.516$\pm$0.006
GRADATE [AAAI’23] 0.746$\pm$0.008 0.686$\pm$0.005 0.790$\pm$0.014 0.705$\pm$0.013 0.805$\pm$0.007 0.719$\pm$0.005 0.574$\pm$0.005 0.544$\pm$0.006
VGOD [ICDE’23] 0.745$\pm$0.002 0.679$\pm$0.003 0.791$\pm$0.005 0.703$\pm$0.003 0.797$\pm$0.001 0.707$\pm$0.006 0.571$\pm$0.006 0.531$\pm$0.008
GAE DOMINANT [arXiv’16] 0.681$\pm$0.013 0.625$\pm$0.011 0.699$\pm$0.019 0.625$\pm$0.011 0.694$\pm$0.007 0.605$\pm$0.005 0.539$\pm$0.003 0.509$\pm$0.003
GCNAE [SDM’19] 0.683$\pm$0.005 0.633$\pm$0.006 0.701$\pm$0.017 0.631$\pm$0.019 0.679$\pm$0.010 0.591$\pm$0.012 0.541$\pm$0.009 0.511$\pm$0.006
AnomalyDAE [ICASSP’20] 0.679$\pm$0.009 0.622$\pm$0.004 0.697$\pm$0.023 0.626$\pm$0.018 0.708$\pm$0.018 0.610$\pm$0.024 0.545$\pm$0.014 0.515$\pm$0.007
AdONE [WSDM’20] 0.685$\pm$0.009 0.624$\pm$0.004 0.707$\pm$0.020 0.637$\pm$0.018 0.721$\pm$0.015 0.634$\pm$0.014 0.556$\pm$0.011 0.526$\pm$0.018
GAD-NR [WSDM’24] 0.749$\pm$0.002 0.691$\pm$0.003 0.789$\pm$0.013 0.710$\pm$0.015 0.829$\pm$0.001 0.721$\pm$0.006 0.579$\pm$0.002 0.550$\pm$0.008
ADA-GAD [AAAI’24] 0.742$\pm$0.009 0.702$\pm$0.010 0.795$\pm$0.003 0.716$\pm$0.004 0.841$\pm$0.016 0.732$\pm$0.014 0.582$\pm$0.003 0.549$\pm$0.006
GADAM [ICLR’24] 0.751$\pm$0.001 0.702$\pm$0.003 0.796$\pm$0.009 0.713$\pm$0.011 0.838$\pm$0.003 0.728$\pm$0.001 0.580$\pm$0.005 0.554$\pm$0.005
MV AnomMAN [IS’23] 0.750$\pm$0.008 0.703$\pm$0.006 0.792$\pm$0.006 0.714$\pm$0.007 0.839$\pm$0.007 0.728$\pm$0.005 0.577$\pm$0.0011 0.552$\pm$0.0010
DualGAD [IS’23] 0.752$\pm$0.005 0.700$\pm$0.008 0.789$\pm$0.013 0.714$\pm$0.014 0.835$\pm$0.008 0.727$\pm$0.004 0.578$\pm$0.006 0.550$\pm$0.009
UMGAD 0.780$\pm$0.004 0.728$\pm$0.007 0.831$\pm$0.003 0.744$\pm$0.0006 0.886$\pm$0.004 0.793$\pm$0.002 0.612$\pm$0.002 0.600$\pm$0.004
Improvement $3.59\%\uparrow$ $3.56\%\uparrow$ $4.40\%\uparrow$ $3.91\%\uparrow$ $5.35\%\uparrow$ $8.33\%\uparrow$ $4.24\%\uparrow$ $3.53\%\uparrow$
V-F Performance Comparison in the Ground Truth Leakage Scenario (RQ5)
In this experiment, following [46], we determine the anomaly score thresholds for all methods on each dataset based on the number of anomalies. The experimental results are presented in Table IV. As shown in Table IV, even when ground truth information is incorporated, our model still achieves the best performance. This demonstrates that UMGAD effectively captures anomaly information and identifies anomalous nodes with greater accuracy compared to other methods.
For traditional GAD methods, Radar utilizes residual networks to learn node attributes and capture anomalies in a simple and effective manner. However, it struggles to capture anomalous features when dealing with complex network structures. Among specifically designed MPI methods, TAM outperforms ComGA and RAND by learning custom node embeddings for anomaly detection. It maximizes the local affinity between nodes and their neighbors, while minimizing the influence of non-homogeneous edges linking abnormal and normal nodes.
CL-based GAD methods, such as GRADATE, VGOD, and ARISE, stand out as the most effective approaches. All three methods emphasize the importance of anomalous structural subgraphs in the GAD problem, leveraging subgraph sampling strategies to learn subgraph features and capture structural anomalies that are difficult for other methods to detect. In the realm of GAE-based GAD methods, most focus on improving the detection of anomalous nodes by learning better features during the reconstruction of attribute and structural information. The four most effective and representative methods in this category are GADAM, AnomMAN, GAD-NR, and ADA-GAD. ADA-GAD enhances anomalous node feature learning through graph augmentation, GAD-NR incorporates neighborhood reconstruction into the GAD task, and GADAM improves anomaly detection performance through enhanced GNN message passing. However, these methods fall short when addressing the GAD problem on multiplex heterogeneous graphs, leading to slightly lower performance compared to our UMGAD. AnomMAN distinguishes itself by targeting multiplex heterogeneous graphs constructed from ordinary heterogeneous graphs, which aligns with the focus of our UMGAD. However, its performance is limited due to its inability to effectively capture subgraph structural anomalies.
(a) Running time of each epoch
(b) Running time of total model
(c) Training Loss
@@ V-F Performance Comparison in the Ground Truth Leakage Scenario (RQ5)
V-G Model Efficiency Analysis (RQ6)
To evaluate the time efficiency of UMGAD, we compare its single-epoch and total runtime across four datasets against the four top-performing SOTA models: ADA-GAD, TAM, GADAM, and AnomMAN. As shown in Fig. 6, although UMGAD is not the fastest on the Amazon dataset in terms of single-epoch and total runtime, it remains highly competitive with these leading approaches. Significantly, on the YelpChi dataset, UMGAD achieves the shortest total runtime among all baselines, highlighting its superior efficiency. Additionally, as depicted in Fig. 6(c), UMGAD demonstrates rapid convergence during training, reaching optimal performance with fewer epochs. This combination of efficiency and robust anomaly detection performance underscores UMGAD’s practicality and scalability, making it a probably choice for large-scale, time-sensitive graph analysis tasks.

### v4

@@ V-B Anomaly Score Threshold Selection for Unsupervised Anomaly Detection (RQ1)
The results shown in Fig. 2 illustrate the trend of ranked node anomaly scores for our UMGAD and the best-performing baselines: ADA-GAD, TAM, GADAM, and AnomMAN on four small-scale datasets, and ADA-GAD, GRADATE, GADAM, and DualGAD on two large-scale datasets. Compared to other baselines, the curve for our UMGAD converges more quickly, stabilizing closer to the actual number of anomalies in the datasets. UMGAD effectively distinguishes anomaly scores between anomalous and normal nodes, enabling more accurate anomaly identification. A robust anomaly detection model should separate anomalous nodes from normal ones using anomaly scores. When these scores are ranked in descending order, a clear inflection point should appear, where the scores stabilize. Nodes with higher scores are likely anomalous, while those with lower, more uniform scores are normal. We use the anomaly score at this inflection point as the threshold, and the number of anomalous nodes identified at this threshold closely matches the actual number of anomalies, as confirmed by ground truth.
Our proposed anomaly threshold selection strategy is the first truly unsupervised method. It does not require any ground-truth information from the test set, making it a fair and objective model evaluation approach. Therefore, the experimental results in Table II are objective. Although the results in Table V are generally better than those in Table II, they are not objective, as ground-truth information from the test set (e.g., the number of anomalous nodes) is used when selecting the anomaly score threshold.
@@ V-C Performance Comparison in the Real Unsupervised Scenario (RQ2)
Next, we evaluate the performance of our UMGAD and all baselines in real unsupervised scenarios. The experimental results are shown in Table II. The best results are highlighted in bold, and the second-best results are underlined. As we can see, our UMGAD obtains the optimal performance, significantly achieves 13.48%, and 11.68% average improvement in terms of AUC and Macro-F1 across four datasets. We can draw the following conclusions: First, according to our proposed threshold selection strategy and the node anomaly score curves in Fig. 2, the number of detected anomalies corresponding to our selected thresholds is the closest to the real situation compared to other methods. Second, UMGAD considers the multi-relational correlation between nodes, and collaboratively combines GAEs with various masking mechanisms across different types of relations, effectively capturing diverse anomaly signals in both original and augmented views. Therefore, the anomaly scores of anomalous nodes can be better distinguished from those of normal nodes.
V-D Performance Comparison in large-sacle graphs (RQ3)
TABLE III: Performance comparison of models on large-scale graph datasets in the real unsupervised scenario.
Method DG-Fin T-Social
AUC Macro-F1 AUC Macro-F1
MPI ComGA 0.580$\pm$0.007 0.501$\pm$0.009 0.549$\pm$0.006 0.489$\pm$0.005
RAND 0.584$\pm$0.009 0.503$\pm$0.011 0.561$\pm$0.005 0.498$\pm$0.002
CL PREM 0.626$\pm$0.007 0.510$\pm$0.009 0.590$\pm$0.003 0.502$\pm$0.005
GRADATE 0.636$\pm$0.010 0.516$\pm$0.011 0.606$\pm$0.009 0.509$\pm$0.007
VGOD 0.632$\pm$0.008 0.511$\pm$0.010 0.597$\pm$0.012 0.503$\pm$0.011
GAE ADA-GAD 0.662$\pm$0.004 0.530$\pm$0.003 0.653$\pm$0.004 0.515$\pm$0.004
GADAM 0.651$\pm$0.009 0.522$\pm$0.011 0.638$\pm$0.011 0.600$\pm$0.010
MV DualGAD 0.663$\pm$0.005 0.532$\pm$0.004 0.648$\pm$0.003 0.510$\pm$0.004
UMGAD 0.733$\pm$0.009 0.598$\pm$0.005 0.712$\pm$0.003 0.558$\pm$0.002
Improvement $10.52\%\uparrow$ $12.52\%\uparrow$ $9.04\%\uparrow$ $8.49\%\uparrow$
To evaluate UMGAD on large-scale datasets DG-Fin and T-Social, we compare it with state-of-the-art methods that avoid out-of-memory (OOM) issues. Only ComGA, RAND, PREM, GRADATE, VGOD, ADA-GAD, GADAM, and DualGAD successfully run. As shown in Table III, UMGAD outperforms all models in AUC and Macro-F1, with average improvements of 9.78% and 10.51%, respectively, over the best baseline. DualGAD performs second, followed by ADA-GAD and GADAM. Specifically, in DGraph-Fin, only 32.2% of users have engaged in borrowing behavior, with true anomalies making up just 1.3%, and anomalies overall accounting for only 0.4%. This extreme imbalance makes detection challenging. Despite this, our method consistently outperforms other SOTA methods in both effectiveness and efficiency, as shown in Table III and Fig. 7. These findings highlight UMGAD’s exceptional efficiency in handling large-scale graphs.
TABLE IV: The comparison of UMGAD and its variants on AUC and macro-F1 (F1 for short) metrics.
@@ V-D Performance Comparison in large-sacle graphs (RQ3)
w/o M 0.712 0.669 0.783 0.689 0.829 0.666 0.865 0.556
w/o O 0.720 0.682 0.788 0.696 0.835 0.678 0.874 0.565
w/o A 0.714 0.670 0.782 0.691 0.827 0.670 0.869 0.558
w/o NA 0.731 0.690 0.795 0.706 0.842 0.686 0.878 0.575
w/o SA 0.726 0.685 0.788 0.701 0.839 0.682 0.875 0.567
w/o DCL 0.748 0.702 0.806 0.718 0.848 0.705 0.883 0.577
UMGAD 0.770 0.722 0.825 0.740 0.878 0.729 0.908 0.605
V-E Ablation Study (RQ4)
@@
Table IV presents the results of model variants on four datasets in real unsupervised scenarios. All five variants underperform UMGAD in terms of AUC and Macro-F1, underscoring the necessity and effectiveness of the proposed components. Among them, w/o M performs the worst, highlighting the importance of the graph-masked autoencoder (GMAE) in GAD tasks. Ordinary GAE struggles to accurately reconstruct the attributes and structural features of anomalous nodes, leading to suboptimal anomaly detection performance. w/o O outperforms w/o M and w/o A, emphasizing the critical role of the original multiplex heterogeneous graph. Relying solely on the augmented graph risks losing crucial anomaly information. The results for w/o NA and w/o SA further demonstrate that the augmented-view GMAE is effective in capturing anomalous information. Changes to node attributes and structural features in the augmented graph help mitigate the homogeneity trap common in GNN-based autoencoders. Finally, while w/o CDL performs the best among all variants, it still falls short of UMGAD, confirming the significant contribution of dual-view contrastive learning in optimizing node representations.
@@
Figure 4: Effect of masking ratio and subgraph size. X-axis represents the mask ratio, while the legend denotes subgraph size.
V-F Parameter Sensitivity Analysis (RQ5)
V-F1 The effect of hyperparameters $\lambda$, $\mu$ and $\Theta$ in the final loss function
$\lambda$, $\mu$ and $\Theta$ correspond to the importance parameters of the two augmented views and the dual-view contrastive learning module, respectively. Considering that incorporating node attribute-level and subgraph-level augmented views on top of the original view can improve model performance, the results of the parameter sensitivity experiments are shown in Fig. 3. Our UMGAD performs optimally when $\lambda$ and $\mu$ are set to 0.3 and 0.3, 0.3 and 0.4, 0.4 and 0.4, and 0.4 and 0.5 on the four small-scale datasets, respectively. Additionally, our UMGAD consistently performs well with $\Theta$ set to 0.1 across all datasets.
V-F2 The effect of masking ratio $r_{m}$ and masking subgraph size $|\mathcal{V}_{m}|$
@@ V-F1 The effect of hyperparameters $\lambda$, $\mu$ and $\Theta$ in the final loss function
Figure 5: The effect of weights $\alpha$ and $\beta$ in four datasets. AUC values for Retail and Alibaba are shown on the left Y-axis, while those for Amazon and YelpChi are on the right Y-axis.
V-F3 The effect of weights $\alpha$ and $\beta$ that balance the importance of attribute and structure reconstruction
We conduct a sensitivity analysis of $\alpha$ and $\beta$, which balance attribute and structure reconstructions in the original and augmented views. As shown in Fig. 5, the model’s performance declines sharply for extreme values of $\alpha$ and $\beta$ $(<0.2\text{ or }>0.8)$, highlighting the importance of maintaining a balance. The AUC value first increases and then decreases as $\alpha$ and $\beta$ grow, with optimal performance of our proposed UMGAD achieved when $\alpha$ and $\beta$ are generally selected from {0.4, 0.5, 0.6} and {0.3, 0.4, 0.5} respectively, across four different datasets.
V-F4 The trade-off between efficiency and accuracy
We evaluate the trade-off between efficiency and accuracy for our UMGAD by analyzing its runtime and AUC scores across different variants. Specifically, Att excludes structural reconstruction, retaining only attribute reconstruction in both the original and augmented views of graphs. Str removes attribute reconstruction, focusing solely on structural reconstruction. Sub only includes the subgraph reconstruction mechanism in both views. For Att and Str, we inject attribute or structural anomalies, respectively, to simulate scenarios where the anomalies are either attribute- or structure-based. Experimental results, shown in Fig. 6, reveal that by selectively pruning the model for different anomaly types, we can enhance efficiency without sacrificing significant performance, adapting to various anomaly scenarios.
(a) Retail
(b) Alibaba
Figure 6: The trade-off between accuracy and efficiency on injected anomaly datasets Retail and Alibaba.
TABLE V: Performance comparison of all models on four datasets in the scenario of anomaly score threshold selection with ground truth leakage.
@@ V-F4 The trade-off between efficiency and accuracy
Trad. Radar [IJCAI’17] 0.678$\pm$0.003 0.570$\pm$0.001 0.695$\pm$0.011 0.621$\pm$0.009 0.659$\pm$0.003 0.543$\pm$0.002 0.515$\pm$0.003 0.402$\pm$0.005
MPI ComGA [WSDM’22] 0.702$\pm$0.009 0.648$\pm$0.003 0.661$\pm$0.008 0.552$\pm$0.002 0.715$\pm$0.004 0.602$\pm$0.007 0.543$\pm$0.011 0.397$\pm$0.006
RAND [ICDM’23] 0.726$\pm$0.006 0.678$\pm$0.007 0.784$\pm$0.008 0.677$\pm$0.004 0.730$\pm$0.013 0.615$\pm$0.006 0.554$\pm$0.007 0.404$\pm$0.008
TAM [NeurIPS’24] 0.753$\pm$0.001 0.703$\pm$0.001 0.790$\pm$0.001 0.715$\pm$0.001 0.848$\pm$0.004 0.703$\pm$0.001 0.582$\pm$0.001 0.425$\pm$0.001
CL CoLA [TNNLS’21] 0.629$\pm$0.009 0.559$\pm$0.006 0.647$\pm$0.014 0.553$\pm$0.003 0.661$\pm$0.006 0.554$\pm$0.002 0.792$\pm$0.010 0.519$\pm$0.009
ANEMONE [CIKM’21] 0.689$\pm$0.004 0.631$\pm$0.004 0.710$\pm$0.007 0.632$\pm$0.009 0.704$\pm$0.002 0.584$\pm$0.005 0.807$\pm$0.006 0.531$\pm$0.003
Sub-CR [IJCAI’22] 0.704$\pm$0.009 0.635$\pm$0.008 0.728$\pm$0.006 0.648$\pm$0.008 0.675$\pm$0.018 0.568$\pm$0.003 0.810$\pm$0.009 0.539$\pm$0.008
ARISE [TNNLS’23] 0.740$\pm$0.018 0.683$\pm$0.005 0.774$\pm$0.006 0.669$\pm$0.003 0.768$\pm$0.006 0.645$\pm$0.002 0.815$\pm$0.001 0.535$\pm$0.007
SL-GAD [TKDE’21] 0.725$\pm$0.003 0.667$\pm$0.007 0.778$\pm$0.005 0.672$\pm$0.003 0.725$\pm$0.009 0.610$\pm$0.010 0.825$\pm$0.013 0.540$\pm$0.005
PREM [ICDM’23] 0.724$\pm$0.001 0.675$\pm$0.003 0.782$\pm$0.011 0.670$\pm$0.009 0.735$\pm$0.007 0.619$\pm$0.006 0.787$\pm$0.003 0.513$\pm$0.007
GCCAD [TKDE’22] 0.731$\pm$0.003 0.677$\pm$0.005 0.788$\pm$0.002 0.689$\pm$0.003 0.733$\pm$0.001 0.617$\pm$0.005 0.796$\pm$0.015 0.516$\pm$0.006
GRADATE [AAAI’23] 0.746$\pm$0.008 0.686$\pm$0.005 0.790$\pm$0.014 0.705$\pm$0.013 0.805$\pm$0.007 0.685$\pm$0.005 0.780$\pm$0.005 0.514$\pm$0.006
VGOD [ICDE’23] 0.745$\pm$0.002 0.679$\pm$0.003 0.791$\pm$0.005 0.703$\pm$0.003 0.797$\pm$0.001 0.677$\pm$0.006 0.771$\pm$0.006 0.512$\pm$0.008
GAE DOMINANT [arXiv’16] 0.681$\pm$0.013 0.625$\pm$0.011 0.699$\pm$0.019 0.615$\pm$0.011 0.694$\pm$0.007 0.598$\pm$0.005 0.539$\pm$0.003 0.389$\pm$0.003
GCNAE [SDM’19] 0.683$\pm$0.005 0.633$\pm$0.006 0.701$\pm$0.017 0.631$\pm$0.019 0.679$\pm$0.010 0.591$\pm$0.012 0.541$\pm$0.009 0.401$\pm$0.006
AnomalyDAE [ICASSP’20] 0.679$\pm$0.009 0.622$\pm$0.004 0.697$\pm$0.023 0.626$\pm$0.018 0.708$\pm$0.018 0.603$\pm$0.024 0.583$\pm$0.035 0.419$\pm$0.007
AdONE [WSDM’20] 0.685$\pm$0.009 0.624$\pm$0.004 0.707$\pm$0.020 0.637$\pm$0.018 0.721$\pm$0.015 0.616$\pm$0.014 0.686$\pm$0.011 0.506$\pm$0.018
GAD-NR [WSDM’24] 0.749$\pm$0.002 0.691$\pm$0.003 0.789$\pm$0.013 0.710$\pm$0.015 0.829$\pm$0.001 0.703$\pm$0.006 0.801$\pm$0.002 0.530$\pm$0.008
ADA-GAD [AAAI’24] 0.742$\pm$0.009 0.702$\pm$0.010 0.795$\pm$0.003 0.716$\pm$0.004 0.837$\pm$0.016 0.706$\pm$0.014 0.832$\pm$0.003 0.552$\pm$0.006
GADAM [ICLR’24] 0.751$\pm$0.001 0.702$\pm$0.003 0.796$\pm$0.009 0.713$\pm$0.011 0.832$\pm$0.003 0.704$\pm$0.001 0.821$\pm$0.005 0.549$\pm$0.005
MV AnomMAN [IS’23] 0.750$\pm$0.008 0.703$\pm$0.006 0.792$\pm$0.006 0.714$\pm$0.007 0.841$\pm$0.007 0.711$\pm$0.005 0.837$\pm$0.011 0.554$\pm$0.010
DualGAD [IS’23] 0.752$\pm$0.005 0.700$\pm$0.008 0.789$\pm$0.013 0.714$\pm$0.014 0.839$\pm$0.008 0.708$\pm$0.004 0.849$\pm$0.006 0.572$\pm$0.014
UMGAD 0.780$\pm$0.004 0.728$\pm$0.007 0.831$\pm$0.003 0.744$\pm$0.006 0.886$\pm$0.004 0.749$\pm$0.002 0.912$\pm$0.002 0.610$\pm$0.004
Improvement $3.59\%\uparrow$ $3.56\%\uparrow$ $4.40\%\uparrow$ $3.91\%\uparrow$ $4.48\%\uparrow$ $8.33\%\uparrow$ $7.42\%\uparrow$ $6.64\%\uparrow$
(a) Running time of each epoch.
(b) Running time of total model.
(c) Training loss on YelpChi.
@@
V-G Performance Comparison in the Ground Truth Leakage Scenario (RQ6)
In this experiment, following [49], we determine the anomaly score thresholds for all methods on each dataset based on the number of anomalies. The experimental results are presented in Table V. As shown in Table V, even when ground truth information is incorporated, our model still achieves the best performance. This demonstrates that UMGAD effectively captures anomaly information and identifies anomalous nodes with greater accuracy compared to other methods.
For traditional GAD methods, Radar utilizes residual networks to learn node attributes and capture anomalies simply and effectively. However, it struggles to capture anomalous features when dealing with complex network structures. Among specifically designed MPI methods, TAM outperforms ComGA and RAND by learning custom node embeddings. It maximizes the local affinity between nodes and their neighbors, while minimizing the influence of non-homogeneous edges linking abnormal and normal nodes.
CL-based GAD methods, such as GRADATE, VGOD, and ARISE, stand out as the most effective approaches. All three methods emphasize the importance of anomalous structural subgraphs in the GAD problem, leveraging subgraph sampling strategies to learn subgraph features and capture structural anomalies that are difficult for other methods to detect. In the realm of GAE-based GAD methods, most focus on improving the detection of anomalous nodes by learning better features during the reconstruction of attribute and structural information. The four most effective and representative methods in this category are GADAM, AnomMAN, GAD-NR, and ADA-GAD. ADA-GAD enhances anomalous node feature learning through graph augmentation, GAD-NR incorporates neighborhood reconstruction into the GAD task, and GADAM improves anomaly detection performance through enhanced GNN message passing. However, these methods fall short when addressing the GAD problem on multiplex heterogeneous graphs, leading to slightly lower performance compared to our UMGAD. AnomMAN focuses on multi-view graphs derived from ordinary heterogeneous graphs, similar to our UMGAD. However, its performance is limited by its inability to effectively capture subgraph structural anomalies.
V-H Model Efficiency Analysis (RQ7)
To evaluate the time efficiency of UMGAD, we compare its single-epoch and total runtime across Retail, YelpChi, and T-Social datasets against four best-performing SOTA models: GRADATE, GADAM, ADA-GAD, and DualGAN. As shown in Fig. 7, UMGAD is faster than other baselines regarding single-epoch and total runtime across all datasets. Significantly, on the large-scale dataset T-Social, UMGAD achieves the shortest epoch and total runtime, highlighting its superior efficiency. Additionally, as shown in Fig. 7(c), UMGAD demonstrates rapid convergence during training, reaching optimal performance with fewer epochs. This combination of efficiency and robust anomaly detection performance underscores UMGAD’s practicality and scalability, making it a probable choice for large-scale, time-sensitive graph analysis tasks.